OpenStudy (anonymous):

prove that 10x-4x=6x. use the algebraic theorems and axioms

OpenStudy (bahrom7893):

@cwrw238 @AravindG @ganeshie8 ideas?

OpenStudy (loser66):

10x-4x= x(10-4)=x6=6x